The fundamental frequency of a guitar string is 388 Hz. What is the fundamental frequency if the tension in the string is reduced by half?
Answer #1
given , fundamental frequency = n1= 388Hz
    intial tension = T1
final tension = T1/2
final fundamental frquency (n2)= ?
we know that = nα√T
so n2/n1 =√T2/T1 =
n2 = n1√T2/T1 = 388√0.5 =274.3Hz